OpenStudy (anonymous):

Find the area of the circle with the given radius or diameter. Use sqrt = 3.14. d = 13 A = ? sq. units

OpenStudy (anonymous):

??????

OpenStudy (stormfire1):

The area of a circle is \[\pi r^2\] The diameter is \[\frac{1}{2}d\] d = 13, so r = 6.5 \[\pi * (6.5)^2 = ~132.73\]

OpenStudy (stormfire1):

Actually, if you're only using 3.14, the final answer would be 132.665

OpenStudy (anonymous):

See You guys make it so much easier

OpenStudy (stormfire1):

Oh, and I meant to say the *radius* is 1/2d...not awake yet

OpenStudy (anonymous):

What if we change it up and the diameter = 12

OpenStudy (stormfire1):

Then instead of squaring 6.5 (whic is 1/2 of 13), you'd square 6 (which is 1/2 of 12). Just remember that the radius is 1/2 of the diameter and you'll be able to solve these whether you're given a radius or a diameter in the problem.